# A Dimca-Greuel type inequality for foliations

### v2

@@ 5. Applications
In this section, from the proof of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
@@
Let $f(x,y)\in\mathbb{C}\{x,y\}$ an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
@@
Consider the singular holomorphic foliation $\mathcal{F}:\omega=df$. In this situation, $C\in\text{ISO}_{0}(\mathcal{F})$. To continue, we compute
$\text{DG}_{0}(\mathcal{F},C)=R_{0}(\mathcal{F},C)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)}),$
where
$R_{0}(\mathcal{F},C):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(2\nu_{q^{(j)}}(C^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
and $C^{(j+1)}$ denotes the strict transform of $C^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $C^{(N)}$ is the normalization of $C$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in C^{(N)}$ as in the proof of Theorem A.
Since $C^{(N)}$ is smooth, we have $\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})=GSV_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})$ by Remark 4.2, and from Lemma 3.1, we deduce
$\text{DG}_{0}(\mathcal{F},C)=-GSV_{0}(\mathcal{F},C)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right]$
Now, since $\mathcal{F}$ is a curve generalized non-dicritical foliation at $0\in\mathbb{C}^{2}$, and $\text{Sep}_{0}(\mathcal{F})=C$, we have $GSV_{0}(\mathcal{F},C)=0$ by [3, p. 528]. Moreover $\mu_{0}(\mathcal{F},C)=\mu(C)$ and $\tau_{0}(\mathcal{F},C)=\tau(C)$ by Remark 2.1. Hence
$\text{DG}_{0}(\mathcal{F},C)=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right].$
Finally, applying successively [19, Claim 4.4] in each step of the resolution of singularities of $C$, and given that the corresponding value $\mathcal{D}_{min}$ of [19, Claim 4.4] in the $j$-th step of the resolution of $C$ is less or equal to $\alpha_{j}$, we get $\text{DG}_{0}(\mathcal{F},C)<0$ implying
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}$
by Theorem A. ∎
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. If $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$ with $\text{DG}_{p}(\mathcal{F},B_{j})<0$ for all $1\leq j\leq\ell$, then
(28) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}<\frac{4}{3}.$
@@
$\sum_{j=1}^{\ell}\left(3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})\right)\leq-\ell<-\ell+1.$
@@
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})<-\ell+1=3\ell-4\ell+(4-3)$
and
$3\left(\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-\ell+1\right)<4\left(\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})-\ell+1\right),$
@@
$\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
@@
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$. Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We have $GSV_{0}(\mathcal{F},C)=5$, $GSV_{0}(\mathcal{F},C_{1,1})=5$, $\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$, hence $\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B})}=\frac{25}{26}<\frac{4}{3}$.
@@
Question 1: Is the inequality (28) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?

### v3

@@ 5. Applications
From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17]:
@@
Let $f(x,y)\in\mathbb{C}\{x,y\}$ be an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
@@
It is a consequence of Remark 2.1 and applying Theorem A to the foliation $\mathcal{F}:df=0$, because $\mu_{p}(\mathcal{F},C)=\mu(C)$, $\tau_{p}(\mathcal{F},C)=\tau(C)$, and $GSV_{p}(\mathcal{F},C)=0$. ∎
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
(21) $03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0
@@
$\sum_{j=1}^{\ell}\left[3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})+GSV_{p}(\mathcal{F},B_{j})\right]\leq-\ell.$
@@
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4T_{p}(\mathcal{F},\mathcal{B})+\sum_{j=1}^{\ell}GSV_{p}(\mathcal{F},B_{j})\leq-\ell$
and, using (7) together with the GSV-index addition formula ([3, p. 29]), we obtain
$3\left(\mu_{p}(\mathcal{F},\mathcal{B})+\ell-1\right)-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})+2\sum_{i\neq k}i_{p}(B_{i},B_{k})\leq-\ell,$
@@
$3\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})\leq-4\ell+3-2\sum_{i\neq k}i_{p}(B_{i},B_{k}).$
Since $\ell\geq 1$, we get
$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0 .
@@
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$. Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We have $GSV_{0}(\mathcal{F},C)=GSV_{0}(\mathcal{F},C_{1,1})=5$, so
$GSV_{0}(\mathcal{F},\mathcal{B})=GSV_{0}(\mathcal{F},C)+GSV_{0}(\mathcal{F},C_{1,1})-2i_{0}(C,C_{1,1})=5+5-2\cdot 5=0,$
$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$. Hence
$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})=3\cdot 25-4\cdot 26=-29<$ 0 .
@@
Question: Is the inequality (21) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?
Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done.
Funding. This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979 and by the Spanish grant PID2019-105896GB-I00 funded by MCIN/AEI/10.13039/501100011033.
Conflict of interest. The authors of this work declare that they have no conflicts of interest.